the radius of a cylinder is 3.6cm and the height is 12.5. what is the volume. please show work
Answer by nerdybill(7384)

the radius of a cylinder is 3.6cm and the height is 12.5. what is the volume. .
.
Volume = "base" times "height"
.
"base" = (pi)r^2
"base" = (3.14)(3.6^2)
.
Volume = (3.14)(3.6^2) * 12.5
Volume = (3.14)(12.96) * 12.5
Volume = 40.6944 * 12.5
Volume = 508.68 cubic cm